# 94% on CIFAR-10 in 3.29 Seconds on a Single GPU

### v1

@@ 3.1 Network architecture and baseline training
As our baseline, we train using Nesterov SGD at batch size 1048, with a label smoothing rate of 0.2. We use a triangular learning rate schedule which starts at the maximum rate, reaches the maximum at 20% of the way through training, and then decreases to zero. For data augmentation we use random horizontal flipping alongside 2-pixel random translation. For translation we use reflection padding (Zagoruyko & Komodakis, 2016) which we found to be better than zero-padding. Note that what we call 2-pixel random translation is equivalent to padding with 2 pixels and then taking a random 32x32 crop. During evaluation we use horizontal flipping test-time augmentation, where the network is run on both a given test image and its mirror and inferences are made based on the average of the two outputs. With optimized choices of learning rate, momentum, and weight decay, this baseline training configuration yields 94% mean accuracy in 45 epochs taking 18.3 A100-seconds.
@@ 5.2 Does alternating flip generalize?
| Heavy RRC | CC(256, 0.875) | 16 | No | 66.78%n=8 | 66.54%n=28 | 66.58%n=28 |
| Heavy RRC | CC(192, 1.0) | 16 | No | 64.43%n=8 | 64.62%n=28 | 64.63%n=28 |
| Light RRC | CC(256, 0.875) | 16 | No | 59.02%n=4 | 61.84%n=26 | 62.19%n=26 |
| Light RRC | CC(192, 1.0) | 16 | No | 61.79%n=4 | 64.50%n=26 | 64.93%n=26 |
| Heavy RRC | CC(256, 0.875) | 16 | Yes | 67.52%n=8 | 67.65%n=28 | 67.60%n=28 |
| Heavy RRC | CC(192, 1.0) | 16 | Yes | 65.36%n=8 | 65.48%n=28 | 65.51%n=28 |
| Light RRC | CC(256, 0.875) | 16 | Yes | 61.08%n=4 | 62.89%n=26 | 63.08%n=26 |
| Light RRC | CC(192, 1.0) | 16 | Yes | 63.91%n=4 | 65.63%n=26 | 65.87%n=26 |
| Light RRC | CC(192, 1.0) | 20 | Yes | not measured | 65.80%n=16 | 66.02%n=16 |
| Heavy RRC | CC(256, 0.875) | 88 | Yes | 72.34%n=2 | 72.45%n=4 | 72.46%n=4 |

### v2

@@ 3.1 Network architecture and baseline training
As our baseline, we train using Nesterov SGD at batch size 1024, with a label smoothing rate of 0.2. We use a triangular learning rate schedule which starts at the maximum rate, reaches the maximum at 20% of the way through training, and then decreases to zero. For data augmentation we use random horizontal flipping alongside 2-pixel random translation. For translation we use reflection padding (Zagoruyko & Komodakis, 2016) which we found to be better than zero-padding. Note that what we call 2-pixel random translation is equivalent to padding with 2 pixels and then taking a random 32x32 crop. During evaluation we use horizontal flipping test-time augmentation, where the network is run on both a given test image and its mirror and inferences are made based on the average of the two outputs. With optimized choices of learning rate, momentum, and weight decay, this baseline training configuration yields 94% mean accuracy in 45 epochs taking 18.3 A100-seconds.
@@ 5.2 Does alternating flip generalize?
| Heavy RRC | CC(256, 0.875) | 16 | No | 66.78% | 66.54% | 66.58% |
| Heavy RRC | CC(192, 1.0) | 16 | No | 64.43% | 64.62% | 64.63% |
| Light RRC | CC(256, 0.875) | 16 | No | 59.02% | 61.84% | 62.19% |
| Light RRC | CC(192, 1.0) | 16 | No | 61.79% | 64.50% | 64.93% |
| Heavy RRC | CC(256, 0.875) | 16 | Yes | 67.52% | 67.65% | 67.60% |
| Heavy RRC | CC(192, 1.0) | 16 | Yes | 65.36% | 65.48% | 65.51% |
| Light RRC | CC(256, 0.875) | 16 | Yes | 61.08% | 62.89% | 63.08% |
| Light RRC | CC(192, 1.0) | 16 | Yes | 63.91% | 65.63% | 65.87% |
| Light RRC | CC(192, 1.0) | 20 | Yes | not measured | 65.80% | 66.02% |
| Heavy RRC | CC(256, 0.875) | 88 | Yes | 72.34% | 72.45% | 72.46% |